OpenStudy (anonymous):

PLEASE HELP!!!!! Which of the following numbers could be used as a common denominator for adding the fractions 2/63 , -4/35, and 8/2. ? 105 63 315 7

OpenStudy (anonymous):

The number wich is divisible by the three denominators.

OpenStudy (anonymous):

What do you mean

OpenStudy (anonymous):

the last one is 8/21

OpenStudy (anonymous):

Divisible means that if you divide one number by another, you´ll get an exact quotient For example , all even numbers are divisible by 2. So in this case, you can verify that 315/63=5 315/35=9 315/21=15 So 315 is divisible by 63,35,21, therefore can be used as common denominator to add that fractions. 315 is also called the least common multiple of 63,35,21..

OpenStudy (anonymous):

thanks